# biblio.bib

### v2

@@ Abstract
This paper completely determines the non-amenability of the mapping class groups of infinite-type surfaces, the mapping class groups of locally finite infinite graphs of higher ranks, and exhibits a class of mapping class groups of trees or rank-one graphs that are amenable.
Date: June 2025.
@@ 1 Introduction
In [mccarthy1985tits] and [ivanov1986algebraic], McCarthy and Ivanov established independently a Tits alternative for subgroups in the mapping class groups of orientable finite-type surfaces, viz. any subgroup of the mapping class group $\operatorname{\mathcal{MCG}}(S)$ of a finite-type surface $S$ either contains a solvable (actually, abelian) subgroup of finite index, or contains a non-abelian free group $\mathbb{F}_{n}$ on $n\geq 2$ generators. One of the central themes in geometric group theory is determining whether various groups satisfy the Tits alternative.
@@
The answer is negative: first, Funar and Kapoujian exhibited a big mapping class group containing Thompson’s group [funar2004universal, Proposition 2.4] and thus not satisfying the Tits alternative. Other examples of big mapping class groups failing the Tits alternative were given in [lanier2020centers] by Lanier and Loving as well as in [aougab2021isometry] by Aougab, Patel and Vlamis, who showed that many big mapping class groups contain every countable group as discrete subgroup. We remark that, together non-amenability of some Burnside groups [adian1979burnside, adyan1983random] and some finitely presented groups [ol2003non], this implies that big mapping class groups can admit non-amenable closed subgroup containing no non-abelian free group. More generally, Allock showed that most big mapping class groups fail to have the Tits alternative [allcock2021most].
Recall that a topological group is amenable if there exists an invariant probability measure on every compact space on which the group acts continuously. A classical result is that every virtually solvable group is an amenable discrete group, whereas any amenable discrete group cannot contain $\mathbb{F}_{2}$ as a subgroup. Hence, the Tits alternative can also be thought of as a strengthening of the amenable/non-amenable dichotomy of the subgroups in the concerned group, equipped with discrete topology.
Although the Tits alternative fails, we may still try to understand how amenable closed subgroups of big mapping class groups look like. In particular, we may ask if the big mapping class groups are themselves amenable or not.
@@
Although Theorem 1.1 is stated and proved for oriantable surfaces, the proof goes verbatim for non-orientable surfaces.
This is a stronger result than [long2025big] which states that the big mapping class groups are not extremely amenable.
We remark that, as mentioned in [long2025big], there is an alternative geometric proof of this result when $S$ is an infinite-type surface with a non-displaceable subsurface of finite type. In this case, one can construct a continuous action of $\operatorname{\mathcal{MCG}}(S)$ on a separable geodesic $\delta$-hyperbolic space $X$ by isometries such that the action is of general type [domat2022big, horbez2022big]. Since a continuous isometric action of an amenable group on such a space cannot be of general type [long2024diverse, Theorem 9.1.1], this implies that $\operatorname{\mathcal{MCG}}(S)$ cannot be amenable.
Outside of the surface world, a graph analogue of big mapping class groups is the mapping class group of a locally finite infinite graph. It can also be considered as the generalisation of the mapping class group of a finite graph, which is a discrete group isomorphic to $\operatorname{Out}(\mathbb{F}_{n})$, if the fundamental group of the finite graph is $\mathbb{F}_{n}\,$. We note that similarly to mapping class groups of finite-type surfaces, $\operatorname{Out}(\mathbb{F}_{n})$ also satisfies the Tits alternative [bestvina2000tits, bestvina2005tits]. Recently, Domat, Hoganson and Kwak have shown that the pure mapping class groups of finite type graphs also enjoy the Tits alternative, but not those of other graphs [domat2023generating]. As a consequence, the Tits alternative also fails in general for the mapping class group of a locally finite infinite graph. So here, we ask about the amenability of these groups and of their closed subgroups.
@@
When the underlying graph is a tree, the mapping class group is isomorphic as topological groups to the homeomorphism group of its ends [bestvina2025groups, Corollary 4.12].
@@
Together with Proposition 4.2, we can conclude from Theorem 1.4 the result below:
@@
Finally, let us remark that both mapping class groups of infinite-type surfaces and mapping class groups of locally finite infinite graphs are non-Archemidean Polish groups (see [long2025big, Proposition2.1] and [bestvina2025groups, Proposition 4.7]). A complete characterisation of the amenability of these groups, as well as their closed subgroups, in terms of the Ramsey property is given in [moore2013amenability].
@@ Acknowledgement.
The author is grateful to Federica Fanoni for encouraging in writing this paper. The author wants to thank Bruno Duchesne, Federica Fanoni and Matteo Tarocchi for inspiring discussion and helpful comments. The author also thanks Brian Udall for valuable suggestions and other feedback received on the Blooming Beasts conference. The author is sponsored by the ANR grant Tremplin ERC-Starting Grant MAGIC (ANR-TERC-StG-0007).
@@ 2 Preliminaries
Proposition 2.1 (Proposition 4.1, [grigorchuk2017amenability]).
@@
We should remind the reader that, unlike the discrete cases, containing a free group $\mathbb{F}_{2}$ does not imply that the concerned group is not amenable, even if it is a discrete subgroup. More generally, the openness condition in (A1) cannot be dropped. An example is the unitary group $\mathrm{U}(\mathcal{H}_{\mathbb{C}}^{\infty})$ of an infinite dimensional separable complex Hilbert space $\mathcal{H}_{\mathbb{C}}^{\infty}\,$. Equipped with the strong operator topology, the group $\mathrm{U}(\mathcal{H}_{\mathbb{C}}^{\infty})$ is amenable [grigorchuk2017amenability, Proposition 5.2] but contains every countable group as a discrete subgroup [grigorchuk2017amenability, Lemma 5.1].
@@
Throughout this paper, we assume that the surfaces are orientable, connected, and without boundary component. Let $S$ be such a surface. We say that $S$ is of finite type, if the fundamental group $\pi_{1}(S)$ is finitely generated, otherwise $S$ will be of infinite type.
@@
The notion of end space was introduced by [freudenthal1931enden] under the general setting of connected locally compact spaces $X\,$. It is the space given by
@@
The topological group $\operatorname{Homeo}^{+}(S)$ has the compact-open topology, inducing the quotient topology on $\operatorname{\mathcal{MCG}}(S)\,$. One can easily see that the mapping class group $\operatorname{\mathcal{MCG}}(S)$ of a finite-type surface $S$ is discrete. But when the surface $S$ becomes infinite-type, the topology changes dramatically: it is not locally compact nor compactly generated [aramayona2020big, Theorem 4.2].
To be precise, the mapping class group $\operatorname{\mathcal{MCG}}(S)$ of an infinite-type surface $S$ is actually isomorphic as topological group to a closed subgroup of the symmetry group $\mathfrak{S}_{\infty}$ of countably infinite set, equipped with the pointwise convergence topology [long2025big, Proposition 2.1]. In particular, [hernandez2018isomorphisms, bavard2020isomorphisms] implies that the topology on $\operatorname{\mathcal{MCG}}(S)$ is the permutation topology induced by its action on the set of isotopy classes of curves on $S$, i.e. the topology is generated by the pointwise stabilisers of finitely many curves. This makes the topological group $\operatorname{\mathcal{MCG}}(S)$ a separable and completely metrisable space, i.e. a Polish group.
@@
Let $\Gamma$ be a locally finite infinite connected graph. The mapping class group of $\Gamma$ can be thought of as a graph analogue to the mapping class group of a surface, whereas the definition is slightly different.
Recall that $f\colon X\to Y$ is a proper homotopy equivalence if it is proper and there is another proper map $g\colon Y\to X$ such that both $f\circ g$ and $g\circ f$ are properly homotopic to the identity. The mapping class group of $\Gamma$, denoted by $\operatorname{\mathcal{MCG}}(\Gamma)$, is the group of proper homotopy equivalences of $\Gamma$ up to proper homotopy.
Similarly to the mapping class groups of infinite-type surfaces, the group $\operatorname{\mathcal{MCG}}(\Gamma)$ is also equipped with a Polish group topology. This group topology is obtained by the quotient of compact-open topology of inverse pairs of proper homotopy equivalences by proper homotopy. See [bestvina2025groups, §4] for more details.
@@
The mapping class group of the graph $\Gamma$ also acts continuously on the end space $\partial(\Gamma)$ and thus gives rise to a homomorphism $\operatorname{\mathcal{MCG}}(\Gamma)\to\operatorname{Homeo}(\partial(\Gamma))$ and we define he pure mapping class group $\operatorname{\mathcal{PMCG}}(\Gamma)$ to be the kernel of this homomorphism. For more details, see [bestvina2025groups, domat2023coarse].
@@
If $\Gamma$ is a locally finite infinite graph with trivial fundamental group, then it becomes a locally finite tree. In this case, the mapping class group $\operatorname{\mathcal{MCG}}(\Gamma)$ will be isomorphic as topological groups to the homeomorphism group $\operatorname{Homeo}(\partial\Gamma)$ of the end space $\partial\Gamma$, equipped with compact-open topology [bestvina2025groups, Corollary 4.12].
@@ 3 Mapping class groups of infinite-type surfaces
by restricting a mapping class to $\Sigma\,$. Using the definition of the compact-open topology and its quotient topology, it is not difficult to see that $\phi$ is a continuous group homomorphism. Moreover, the group $\operatorname{\mathcal{MCG}}(\Sigma,\partial\Sigma)$ injects in $\operatorname{\mathcal{MCG}}(S)_{(F\cup F^{\prime})}$ since $F\cap\Sigma=\emptyset$, which allows us to deduce that $\phi$ is an epimorphism. But $\operatorname{\mathcal{MCG}}(\Sigma,\partial\Sigma)$ is not amenable, in particular it is a discrete group containing a non-abelian free group of rank $2$ [farb2011primer, Theorem 3.14]. By (A2), the open subgroup $\operatorname{\mathcal{MCG}}(S)_{(F\cup F^{\prime})}$ is not amenable. Now (A1) will yield that $G$ itself is not amenable. ∎
As we have already seen, Theorem 1.1 can also be stated in the following way: all amenable subgroups in the big mapping class group have to be nowhere dense. However, this does not imply that such a group is discrete. For example, we can take the subgroup $G<\operatorname{\mathcal{MCG}}(S)$ generated by the Dehn twists along infinitely many pairwise disjoint curves. It is easy to see that $G$ is isomorphic to $\mathbb{Z}^{\infty}$ and is an abelian group and thus amenable, but having an accumulation point at $\operatorname{Id}\in\operatorname{\mathcal{MCG}}(S)$, so it is not discrete.
@@ 4 Mapping class groups of infinite-type graphs
Let $\Gamma$ be a locally finite infinite-type graph, and let $\operatorname{\mathcal{MCG}}(\Gamma)$ be its mapping class group. Suppose now that $\Gamma$ is not a tree. Recall that a $H\colon C(X,X)\times[0,1]\to C(X,X)$ between two continuous maps $u=H(u,0)\colon X\to X$ and $v=H(u,1)\colon X\to X$ is stationary on $K\subset X$ if the restriction $H(u,t)|_{K}=\operatorname{Id}_{K}$ is the identity map on $K$, for all $t\in[0,1]\,$. Let $K\subset\Gamma$ be a finite subgraph and define $U_{K}$ to be the set of properly homotopy equivalence classes $[f]\in\operatorname{\mathcal{MCG}}(\Gamma)$ with a representative such that
@@
    there is a representative $g$ of $[f]^{-1}\in\operatorname{\mathcal{MCG}}(\Gamma)$ verifying (1) and (2),
@@
    there are proper homotopies $gf\simeq\operatorname{Id}$ and $fg\simeq\operatorname{Id}$ that are stationary on $K$ and preserve complementary components of $K\,$.
In view of [bestvina2025groups, Proposition 4.7], the subgroups $U_{K}$ are open in $\operatorname{\mathcal{MCG}}(\Gamma)$ and form a neighbourhood basis of $\operatorname{Id}\in\operatorname{\mathcal{MCG}}(\Gamma)\,$.
@@
and this homomorphism is continuous since $\operatorname{\mathcal{MCG}}(K^{\prime})$ is discrete by definition and $\ker(\psi)\supset U_{K\cup K^{\prime}}$ is open. We note that the (properly) homotopy classes of homeomorphisms $K^{\prime}\to K^{\prime}$ are also contained in $U_{K}$, which gives us inclusion $\operatorname{\mathcal{MCG}}(K^{\prime})\hookrightarrow U_{K}\,$. This implies that $\psi$ is an epimorphism. However, there exists a natural homomorphism $\operatorname{\mathcal{MCG}}(K^{\prime})\to\operatorname{Out}(\pi_{1}(K^{\prime}))\simeq\operatorname{Out}(\pi_{1}(\mathbb{F}_{n}))$, where $n\geq 2$ is the rank of $K^{\prime}$. But $\operatorname{Out}(\mathbb{F}_{n})$ contains non-abelian free subgroups (see, for example, [bestvina2000tits, §3.4]) as discrete group, so it is not amenable. By (A2), we can conclude that the topological group $U_{K}$ is not amenable as it continuously maps onto $\operatorname{Out}(\mathbb{F}_{n})$ via $\psi\,$. ∎
@@
Indeed, start with a finite subgraph $K^{\prime}\subset\Gamma$ of rank at least $2$. Let $V$ be a compact neighbourhood of $K^{\prime}$ such that no complementary component of $V$ is compact. This $V$ always exists since $\Gamma$ is locally finite, the total number of complementary components of $V$ is finite, and we can take a new $V$ to be the union of $V$ and all its compact complementary components, which remains a compact neighbourhood of $K^{\prime}\,$. Let $K$ be the finite graph $V\setminus K^{\prime}\,$. The $K$ will be a finite subgraph in $\Gamma$ satisfying the assumptions in Lemma 4.1. Thus, $U_{K}$ is not amenable, forcing $\operatorname{\mathcal{MCG}}(\Gamma)$ to be non-amenable in view of (A1).
Suppose now $\Gamma$ is of infinite type. Let $G$ be an arbitrary open subgroup of $\operatorname{\mathcal{MCG}}(\Gamma)\,$. By [bestvina2025groups, Proposition 4.7], there exists a finite subgraph $K_{0}\subset\Gamma$ such that $U_{K_{0}}<G\,$. Now take $K^{\prime}\subset\Gamma\setminus K_{0}$ to be finite non-tree subgraph, which is always feasible since $\Gamma$ is infinite-type. Similarly to the case above, we can construct a compact neighbourhood $V$ of $K^{\prime}\cup K_{0}$ such that $K\coloneqq V\setminus K^{\prime}$ satisfies the assumptions of Lemma 4.1. Note that $U_{K}<U_{K_{0}}<G$ as $K_{0}\subset K\,$. Lemma 4.1 implies that $U_{K}$ is a non-amenable open subgroup in $G$. Hence, we can conclude the non-amenability of $G$ by appealing to (A1). ∎
@@
Note that $\operatorname{\mathcal{MCG}}(T)\simeq\operatorname{Homeo}(\partial T)$ and $\operatorname{Homeo}(\partial T)=\operatorname{\mathcal{MCG}}(\Gamma)/\operatorname{\mathcal{PMCG}}(\Gamma)$. To conclude the desired result, by (A4) and (A5), it suffices to show that the normal closed subgroup $\operatorname{\mathcal{PMCG}}(\Gamma)$ of $\operatorname{\mathcal{MCG}}(\Gamma)$ is amenable. By [bestvina2025groups, Corollary 3.9], we have

### v3

# Non-amenability of mapping class groups of infinite-type surfaces and graphs

@@ Abstract
This paper completely determines the non-amenability of the mapping class groups of infinite-type surfaces, the mapping class groups of locally finite infinite graphs of higher ranks, gives an example of non-amenable stabiliser of a point at infinity of a coarsely bounded generated hyperbolic Polish group, and exhibits a class of mapping class groups of trees or rank-one graphs that are amenable.
Date: January 2026.
@@ 1 Introduction
In [MCC85] and [IVA86], McCarthy and Ivanov independently established a Tits alternative for subgroups in the mapping class groups of orientable finite-type surfaces, viz. any subgroup of the mapping class group $\operatorname{\mathcal{MCG}}(S)$ of a finite-type surface $S$ either contains a solvable (actually, abelian) subgroup of finite index, or contains a non-abelian free group $\mathbb{F}_{n}$ on $n\geq 2$ generators. One of the central themes in geometric group theory is determining whether various groups satisfy the Tits alternative.
@@
The answer is negative: first, Funar and Kapoujian exhibited a big mapping class group containing Thompson’s group [FK04, Proposition 2.4] and thus not satisfying the Tits alternative. Other examples of big mapping class groups failing the Tits alternative were given in [LL20] by Lanier and Loving as well as in [APV21] by Aougab, Patel and Vlamis, who showed that many big mapping class groups contain every countable group as discrete subgroup. We remark that, together non-amenability of some Burnside groups [ADI79, ADI83] and some finitely presented groups [OS03], this implies that big mapping class groups can admit non-amenable closed subgroup containing no non-abelian free group. More generally, Allock showed that most big mapping class groups fail to have the Tits alternative [ALL21].
Recall that a topological group is amenable if there exists an invariant probability measure on every compact space on which the group acts continuously. A classical result is that every virtually solvable group is an amenable discrete group, whereas any amenable discrete group cannot contain $\mathbb{F}_{2}$ as a subgroup. Hence, the Tits alternative can also be thought of as a strengthening of the amenable/non-amenable dichotomy of the subgroups in the concerned group, equipped with discrete topology, i.e. a subgroup is non-amenable if and only if it contains non-abelian free groups.
Although the Tits alternative fails, one may still hope that a similar strengthening of amenable/non-amenable dichotomy in big mapping class group is true for Polish topology. In particular, we may still try to understand how amenable closed subgroups of big mapping class groups look like and to see if the big mapping class groups are themselves amenable or not.
@@
Although Theorem 1.1 is stated and proved for orientable surfaces, the proof goes verbatim for non-orientable surfaces.
This is a stronger result than [LON25] which states that the big mapping class groups are not extremely amenable.
We remark that, as mentioned in [LON25], there is an alternative geometric proof of this result when $S$ is an infinite-type surface with a non-displaceable subsurface of finite type. In this case, one can construct a continuous action of $\operatorname{\mathcal{MCG}}(S)$ on a separable geodesic $\delta$-hyperbolic space $X$ by isometries such that the action is of general type [DD22, HQR22]. Since a continuous isometric action of an amenable group on such a space cannot be of general type [LON24, Theorem 9.1.1], this implies that $\operatorname{\mathcal{MCG}}(S)$ cannot be amenable.
Although geometric group theory is most commonly applied to finitely generated groups, Rosendal [ROS21] observed that the finiteness or compactness condition can be relaxed by using a notion of boundedness for general topological groups, and it has been seen that this allows us to extend the tools of geometric group theory to a larger class of groups, namely coarsely bounded generated, or, for abbreviation, CB generated Polish group. For example, Mann and Rafi [MR23] characterise mapping class groups that are CB generated, and hence have a canonical (up to quasi-isometry) word metric. Moreover, there are CB generated big mapping class groups that is Gromov-hyperbolic [SCH24]. In view of these, in a recent note [CPV21], Chandran, Patel and Vlamis proposed the following problem:
Problem 1.3.
Develop the theory of hyperbolic Polish groups.
This problem is also related to the Tits alternative: it is a classical result from geometric group theory that finitely generated hyperbolic groups satisfies the Tits alternative. In particular, the stabiliser of a point at infinity on the Gromov boundary of a finitely generated hyperbolic group is amenable, see for example [Gd13, Chapitre 8]. However, this is not true in general for CB generated hyperbolic Polish group:
Theorem 1.4.
There exists a CB generated hyperbolic Polish group such that there is a point on its Gromov boundary with a non-amenable stabiliser in it.
The example is given by a big mapping class group and Theorem 1.4 is a consequence of Theorem 1.1.
Outside of the surface world, a graph analogue of big mapping class groups is the mapping class group of a locally finite infinite graph. It can also be considered as the generalisation of the mapping class group of a finite graph, which is a discrete group isomorphic to $\operatorname{Out}(\mathbb{F}_{n})$, if the fundamental group of the finite graph is $\mathbb{F}_{n}\,$. We note that similarly to mapping class groups of finite-type surfaces, $\operatorname{Out}(\mathbb{F}_{n})$ also satisfies the Tits alternative [BFH00, BFH05]. Recently, Domat, Hoganson and Kwak have shown that the pure mapping class groups of finite type graphs also enjoy the Tits alternative, but not those of other graphs [DHK23b]. As a consequence, the Tits alternative also fails in general for the mapping class group of a locally finite infinite graph. So here, we ask about the amenability of these groups and of their closed subgroups.
@@
When the underlying graph is a tree, the mapping class group is isomorphic as topological groups to the homeomorphism group of its ends [AB25, Corollary 4.12].
@@
Together with Proposition 4.2, we can conclude from Theorem 1.6 the following result:
@@
Finally, let us remark that both mapping class groups of infinite-type surfaces and mapping class groups of locally finite infinite graphs are non-Archemidean Polish groups (see [LON25, Proposition2.1] and [AB25, Proposition 4.7]). A complete characterisation of the amenability of these groups, as well as their closed subgroups, in terms of the Ramsey property is given in [MOO13].
@@ Acknowledgement.
The author is grateful to Federica Fanoni for encouraging in writing this paper. The author wants to thank Bruno Duchesne, Federica Fanoni, Matteo Tarocchi and Brian Udall for inspiring discussion and helpful comments. The author also thanks other feedback received on the Blooming Beasts conference and from the anonymous referees. The author is sponsored by the ANR grant Tremplin ERC-Starting Grant MAGIC (ANR-TERC-StG-0007).
@@ 2 Preliminaries
Proposition 2.1 (Proposition 4.1, [Gd17]).
@@
We should remind the reader that, unlike the discrete cases, containing a free group $\mathbb{F}_{2}$ does not imply that the concerned group is not amenable, even if it is a discrete subgroup. More generally, the openness condition in (A1) cannot be removed. An example is the unitary group $\mathrm{U}(\mathcal{H}_{\mathbb{C}}^{\infty})$ of an infinite dimensional separable complex Hilbert space $\mathcal{H}_{\mathbb{C}}^{\infty}\,$. Equipped with the strong operator topology, the group $\mathrm{U}(\mathcal{H}_{\mathbb{C}}^{\infty})$ is amenable [Gd17, Proposition 5.2] but contains every countable group as a discrete subgroup [Gd17, Lemma 5.1].
@@
Throughout this paper, we assume that the surfaces are orientable, connected, and without boundary components. Let $S$ be such a surface. We say that $S$ is of finite type, if the fundamental group $\pi_{1}(S)$ is finitely generated, otherwise $S$ will be of infinite type.
@@
The notion of end space was introduced by [FRE31] under the general setting of connected locally compact spaces XX\,. It is the space given by
@@
The topological group $\operatorname{Homeo}^{+}(S)$ has the compact-open topology, inducing the quotient topology on $\operatorname{\mathcal{MCG}}(S)\,$. One can easily see that the mapping class group $\operatorname{\mathcal{MCG}}(S)$ of a finite-type surface $S$ is discrete. But when the surface $S$ becomes infinite-type, the topology changes dramatically: it is not locally compact nor compactly generated [AV20, Theorem 4.2].
To be precise, the mapping class group $\operatorname{\mathcal{MCG}}(S)$ of an infinite-type surface $S$ is actually isomorphic as topological group to a closed subgroup of the symmetry group $\mathfrak{S}_{\infty}$ of countably infinite set, equipped with the pointwise convergence topology [LON25, Proposition 2.1]. In particular, [HMV18, BDR20] implies that the topology on $\operatorname{\mathcal{MCG}}(S)$ is the permutation topology induced by its action on the set of isotopy classes of curves on $S$, i.e. the topology is generated by the pointwise stabilisers of finitely many curves. This makes the topological group $\operatorname{\mathcal{MCG}}(S)$ a separable and completely metrisable space, i.e. a Polish group.
@@
Let $\Gamma$ be a locally finite, infinite, connected graph. The mapping class group of $\Gamma$ can be thought of as a graph analogue to the mapping class group of a surface, whereas the definition is slightly different.
Recall that $f\colon X\to Y$ is a proper homotopy equivalence if it is proper and there is another proper map $g\colon Y\to X$ such that both $f\circ g$ and $g\circ f$ are properly homotopic to the identity. The mapping class group of $\Gamma$, denoted by $\operatorname{\mathcal{MCG}}(\Gamma)$, is the group of proper homotopy equivalences of $\Gamma$ up to proper homotopy. We remark that this group should not be confused with the group of proper homotopy equivalences of $\Gamma$, which is a proper extension of the mapping class group, nor should it be confused with the group of homotopy equivalences which does not require the properness condition, see [AB25] for further details.
Similarly to the mapping class groups of infinite-type surfaces, the group $\operatorname{\mathcal{MCG}}(\Gamma)$ is also equipped with a Polish group topology. This group topology is obtained by the quotient of compact-open topology of inverse pairs of proper homotopy equivalences by proper homotopy. See [AB25, §4] for more details.
@@
The mapping class group of the graph $\Gamma$ also acts continuously on the end space $\partial(\Gamma)$ and thus gives rise to a homomorphism $\operatorname{\mathcal{MCG}}(\Gamma)\to\operatorname{Homeo}(\partial(\Gamma))$ and we define the pure mapping class group $\operatorname{\mathcal{PMCG}}(\Gamma)$ to be the kernel of this homomorphism. For more details, see [AB25, DHK23a].
@@
If $\Gamma$ is a locally finite infinite graph with trivial fundamental group, then it becomes a locally finite tree. In this case, the mapping class group $\operatorname{\mathcal{MCG}}(\Gamma)$ will be isomorphic as topological groups to the homeomorphism group $\operatorname{Homeo}(\partial\Gamma)$ of the end space $\partial\Gamma$, equipped with compact-open topology [AB25, Corollary 4.12].
@@ 3 Mapping class groups of infinite-type surfaces
by restricting a mapping class to $\Sigma\,$. Using the definition of the compact-open topology and its quotient topology, it is not difficult to see that $\phi$ is a continuous group homomorphism. Moreover, the group $\operatorname{\mathcal{MCG}}(\Sigma,\partial\Sigma)$ injects in $\operatorname{\mathcal{MCG}}(S)_{(F\cup F^{\prime})}$ since $F\cap\Sigma=\emptyset$, which allows us to deduce that $\phi$ is an epimorphism. But $\operatorname{\mathcal{MCG}}(\Sigma,\partial\Sigma)$ is not amenable, in particular it is a discrete group containing a non-abelian free group of rank $2$ [FM11, Theorem 3.14]. By (A2), the open subgroup $\operatorname{\mathcal{MCG}}(S)_{(F\cup F^{\prime})}$ is not amenable. Now (A1) will yield that $G$ itself is not amenable. ∎
As we have already seen, Theorem 1.1 can also be stated in the following way: all amenable subgroups in the big mapping class group have to be nowhere dense. However, this does not imply that such a group is discrete. For example, we can take the subgroup $G<\operatorname{\mathcal{MCG}}(S)$ generated by the Dehn twists along infinitely many pairwise disjoint curves. It is easy to see that $G$ is isomorphic to $\mathbb{Z}^{\infty}$ and is an abelian group, so it is amenable, but has an accumulation point at $\operatorname{Id}\in\operatorname{\mathcal{MCG}}(S)$, so it is not discrete.
Proof of Theorem 1.4.
Let $S\coloneqq\mathbf{S}^{2}\setminus(\mathcal{C}\cup\{p\})$, where $\mathbf{S}^{2}$ is the $2$-dimensional sphere, $\mathcal{C}$ is a Cantor set and $p\in\mathbf{S}^{2}\setminus\mathcal{C}$ is an isolated puncture. It has been already seen in [SCH24] that $\operatorname{\mathcal{MCG}}(S)$ is CB generated hyperbolic Polish group with infinite diameter and is quasi-isometric to the loop graph $L(S;p)$. Consider a non-compact finite type subsurface $V\subset S$ with $p$ being an end of $V$. Let $\varphi\in\operatorname{\mathcal{MCG}}(S)$ be such that $\varphi|_{V}$ is pseudo-Anosov and $\varphi|_{S\setminus V}$ is identity. Since $\varphi$ is loxodromic on $L(S;p)$ by [RAS21, Theorem 1.1], it is also loxodromic on $\operatorname{\mathcal{MCG}}(S)$. Let $\xi$ be a point on the Gromov boundary of $\operatorname{\mathcal{MCG}}(S)$ fixed by $\varphi$. By Lemma 3.1, we can find an open subgroup $H<\operatorname{\mathcal{MCG}}(S)$ fixing $V$ pointwise. We can check that $H$ commutes with $\varphi$, so it is contained in the stabiliser of $\xi$, i.e. the stabiliser of $\xi$ is open and thus not amenable by Theorem 1.1. ∎
@@ 4 Mapping class groups of infinite-type graphs
Let $\Gamma$ be a locally finite infinite-type graph, and let $\operatorname{\mathcal{MCG}}(\Gamma)$ be its mapping class group. Suppose now that $\Gamma$ is not a tree. Recall that a homotopy $H\colon C(X,X)\times[0,1]\to C(X,X)$ between two continuous maps $u=H(u,0)\colon X\to X$ and $v=H(u,1)\colon X\to X$ is stationary on $K\subset X$ if the restriction $H(u,t)|_{K}=\operatorname{Id}_{K}$ is the identity map on $K$, for all $t\in[0,1]\,$. Let $K\subset\Gamma$ be a finite subgraph and define $U_{K}$ to be the set of properly homotopy equivalence classes $f]\in\operatorname{\mathcal{MCG}}(\Gamma)$ with a representative such that
@@
    there is a representative $g$ of $f]^{-1}\in\operatorname{\mathcal{MCG}}(\Gamma)$ verifying (1) and (2),
@@
    there are proper homotopies $gf\simeq\operatorname{Id}$ and $fg\simeq\operatorname{Id}$ that are stationary on $K$ and preserve complementary components of KK\,.
In view of [AB25, Proposition 4.7], the subgroups $U_{K}$ are open in $\operatorname{\mathcal{MCG}}(\Gamma)$ and form a neighbourhood basis of $\operatorname{Id}\in\operatorname{\mathcal{MCG}}(\Gamma)\,$.
@@
and this homomorphism is continuous since $\operatorname{\mathcal{MCG}}(K^{\prime})$ is discrete by definition and $\ker(\psi)\supset U_{K\cup K^{\prime}}$ is open. We note that the (properly) homotopy classes of homeomorphisms $K^{\prime}\to K^{\prime}$ are also contained in $U_{K}$, which gives us inclusion $\operatorname{\mathcal{MCG}}(K^{\prime})\hookrightarrow U_{K}\,$. This implies that $\psi$ is an epimorphism. However, there exists a natural homomorphism $\operatorname{\mathcal{MCG}}(K^{\prime})\to\operatorname{Out}(\pi_{1}(K^{\prime}))\simeq\operatorname{Out}(\pi_{1}(\mathbb{F}_{n}))$, where $n\geq 2$ is the rank of $K^{\prime}$. But $\operatorname{Out}(\mathbb{F}_{n})$ contains non-abelian free subgroups (see, for example, [BFH00, §3.4]) as a discrete group, so it is not amenable. By (A2), we can conclude that the topological group $U_{K}$ is not amenable as it continuously maps onto $\operatorname{Out}(\mathbb{F}_{n})$ via $\psi\,$. ∎
@@
Indeed, start with a finite subgraph $K^{\prime}\subset\Gamma$ of rank at least $2$. Let $V$ be a compact neighbourhood of $K^{\prime}$ such that no complementary component of $V$ is compact. This $V$ always exists since $\Gamma$ is locally finite, the total number of complementary components of $V$ is finite, and we can take a new $V$ to be the union of $V$ and all its compact complementary components, which will remain a compact neighbourhood of $K^{\prime}\,$. Let $K$ be the finite graph $V\setminus K^{\prime}\,$. The $K$ will be a finite subgraph in $\Gamma$ satisfying the assumptions in Lemma 4.1. Thus, $U_{K}$ is not amenable, forcing $\operatorname{\mathcal{MCG}}(\Gamma)$ to be non-amenable in view of (A1).
Suppose now $\Gamma$ is of infinite type. Let $G$ be an arbitrary open subgroup of $\operatorname{\mathcal{MCG}}(\Gamma)\,$. By [AB25, Proposition 4.7], there exists a finite subgraph $K_{0}\subset\Gamma$ such that $U_{K_{0}}<G\,$. Now take $K^{\prime}\subset\Gamma\setminus K_{0}$ to be finite non-tree subgraph, which is always feasible since $\Gamma$ is infinite-type. Similarly to the case above, we can construct a compact neighbourhood $V$ of $K^{\prime}\cup K_{0}$ such that $K\coloneqq V\setminus K^{\prime}$ satisfies the assumptions of Lemma 4.1. Note that $U_{K}<U_{K_{0}}<G$ as $K_{0}\subset K\,$. Lemma 4.1 implies that $U_{K}$ is a non-amenable open subgroup in $G$. Hence, we can conclude the non-amenability of $G$ by appealing to (A1). ∎
@@
Note that $\operatorname{\mathcal{MCG}}(T)\simeq\operatorname{Homeo}(\partial T)$ and $\operatorname{Homeo}(\partial T)=\operatorname{\mathcal{MCG}}(\Gamma)/\operatorname{\mathcal{PMCG}}(\Gamma)$. To conclude the desired result, by (A4) and (A5), it suffices to show that the normal closed subgroup $\operatorname{\mathcal{PMCG}}(\Gamma)$ of $\operatorname{\mathcal{MCG}}(\Gamma)$ is amenable. By [AB25, Corollary 3.9], we have